OpenStudy (anonymous):

if the eigenspace is reduced to the matrix: |10| |01| what would the basis be or v1? wouldn't it just be v1=[0,0]

OpenStudy (anonymous):

yep pretty much since \[\left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right]\left(\begin{matrix}x _{1} \\ x _{2}\end{matrix}\right)=\left(\begin{matrix}0 \\ 0\end{matrix}\right)\]

OpenStudy (anonymous):

so here x1+0x2=0 and 0x1+x2=0 we typically just use one equation to determine its eigenvector

OpenStudy (anonymous):

yeah that what i thought, but i got it wrong in my quiz

OpenStudy (anonymous):

but thanks for clearing it up

OpenStudy (anonymous):

so its wrong?

OpenStudy (anonymous):

the question was: Define T: R2 right arrow R2 by T(x) = Ax, where A = 1 −6 2 −6 . Find a basis B = {v1, v2 } for R2 with the property that [T]B is diagonal.

OpenStudy (anonymous):

i found the eigen vectors to be -1 and -5

OpenStudy (anonymous):

and when i put them into the matrix to get the eigen vectors, i got the reduced matrix as above, |10| |01|

OpenStudy (anonymous):

sorry i mean eigen values to be -1 and -5

OpenStudy (anonymous):

i don't think they are your eigen values..

OpenStudy (anonymous):

you should have eigen values -3 and -2

OpenStudy (anonymous):

do you want to go through this?

OpenStudy (anonymous):

oh right, no i just re did it

OpenStudy (anonymous):

yeah made a mistake with the characteristic equation, thanks for clearing it up!

OpenStudy (anonymous):

yep sweet. just make sure you carefully solve for the determinant. critical step where everyone can get it wrong and hence your solutions will be wrong

OpenStudy (anonymous):

you can do it here if you want so i can see your thought process